# Epistemic diversity across language models mitigates knowledge collapse

### v1

@@ 6. Discussion
In our study of AI ecosystem evolution, we find an optimal degree of diversity greater than zero for the performance of recursively trained models (see Figure 3). The identified optimal degree of diversity in ecosystems of models offers two key takeaways: 1) multiple community models are more successful at preserving the original information than a single model, even when trained on smaller sample sizes; however, 2) the benefits of epistemic diversity peak, but once the number of models becomes too large (and the subsets too small), this capacity decreases again. In the face of AI monoculture, our findings suggest that diversifying models is an essential factor, alongside existing data curation strategies (Xing et al., 2025; Dohmatob et al., 2024b; Feng et al., 2024a; Sorscher et al., 2023; Xie et al., 2023), for preventing collapse in human, AI-shaped knowledge.
In our experiments, conducted with small models and datasets under controlled conditions, increased diversity does mitigate, but does not entirely prevent, model collapse. Given the simplified nature of our setup, it is possible that, in more realistic scenarios, the benefits of diversity could be even more pronounced. In real-world deployments, both models and datasets are far larger, and diversity can arise from additional dimensions of system design, including model architecture, alignment approaches, user interface design, and system messaging. Combining the variations of these dimensions could amplify the benefits of diversity overall. It is also important to consider that real AI ecosystems consist of multiple but similar models. It is often assumed [^6] Of course, we wouldn’t have to assume if information about the training datasets of dominant AI models were made accessible to independent researchers. that these models are trained on the majority of accessible internet (Villalobos et al., 2024; Mozilla, 2024). If this true, then segmentation through reduction, which is the approach we took in this study, is the only approach to training-data diversification for these larg-scale, real-world AI ecosystems. If this is not the case, then diversity could be achieved through expansion by having, say, Gemini train on one portion of the internet while ChatGPT trains on another portion of the internet. Moreover, a more realistic evaluation of community models would use a community-specific, rather than a universal ground-truth, dataset, which would reduce the measured perplexity. Finally, prevailing AI systems today (e.g., OpenAI’s ChatGPT, Google’s Gemini, Anthropic’s Claude) are based on much larger models trained on significantly larger datasets. Prior work indicates a counterintuitive negative relationship between model size and stability when evaluated on a single model (Herel and Mikolov, 2024; Wright et al., 2025; Dohmatob et al., 2024a). Overall, the relationship between these factors and collapse remains insufficiently understood. Examining how the location of the optimal diversity point shifts in more realistic settings and along additional dimensions is an important direction for future research.
Our findings indicate that diversifying AI is essential for long-term knowledge production. This work contributes to a growing body of AI research that challenges the dominance of single, universal models, highlighting the benefits of diversity in contexts such as justice (Jain et al., 2024; Blili-Hamelin et al., 2025; Gebru and Torres, 2024) and safety (Lu and Kleek, 2024). Knowledge collapse is a gradual process that will likely be difficult to detect, especially in early stages. We therefore welcome recent approaches to monitoring (the progress of) epistemic diversity within models and data (Wright et al., 2025; Zhang et al., 2025a) and recommend extending such analyses to the ecosystem level. To not only monitor but also steer the ecosystem of models, a discussion that includes the public is necessary. AI development relies on public resources, not least in the form of training data. Given the high stakes, a discussion among researchers, industry, policymakers, and the public is needed to determine how these resources should be distributed to ensure a just and successful AI ecosystem in the long run.
Appendix A Probability distributions
Figure 6 illustrates the probability distributions of the information space.
Figure 6. Illustration of probability distributions of information space. Because the training data is finite, the empirical distribution, $\hat{p}_{data(t=0)}$, diverges from the true distribution, ${p}_{data}$. The model’s inherent biases (functional expressivity and functional approximation errors) further distort the distribution it represents, which is $p_{\theta(t=0)}$, and the distribution of the data it generates, which is $\hat{p}_{data(t=1)}$.
Appendix B Model progress including performance on original models
@@
Figure 7 shows the perplexity progress of the two model types across ten iteration steps, including the models’ perplexity for the off-the-shelf models.

### v2

@@ 6. Discussion
In our study of AI ecosystem evolution, we find that the optimal level of diversity increases monotonically with the number of self-training iterations (see Figure 3).
This offers a new perspective on generative model performance. Training a single model on as much data as possible is short-sighted: over longer horizons, community models trained on segments of the available data outperform a single model with access to the full dataset. The more often artificial data is recycled for training, the greater the degree of segmentation needed to attain the optimal diversity level. Put simply, the benefits of diversity grow with the number of iterations.
This observed effect is robust across model families, model sizes, dataset scales, and experimental variations involving real data and temperature-based sampling. Moreover, we find that the benefits of diversity also grow as both models and datasets scale. Taken together, our results suggest that differences between models are a key mechanism for mitigating collapse even under more realistic conditions where models are larger, trained on far more data, can be optimized for diverse outputs, and where human-generated data accumulates alongside model-generated data.
Before turning to sociotechnical implications for AI evaluation and governance, we briefly restate the technical aspects from Section 5. To interpret our empirical results, we introduce effective data quality (EDQ), which treats the informativeness of data as a function of both the model distribution and the true distribution [^8] Importantly, even if the EDQ hypothesis were falsified, the empirically observed diversity optimum would remain.. When EDQ is sufficiently high, additional training improves performance; when EDQ is low, further training is harmful and performance declines. This framing has two broad implications for recursively trained models and machine learning more generally. First, gains and losses depend on EDQ regardless of whether the data is human- or model-generated (Wang et al., 2025b; Feng et al., 2024a; Antonello et al., 2021). Second, in low-EDQ regimes, standard scaling laws invert: performance degradation increases with scale, consistent with prior findings (Dohmatob et al., 2024a; Herel and Mikolov, 2024; Wright et al., 2025). Through the lens of EDQ, diversity helps prevent collapse for two reasons: (i) once the low-EDQ regime is reached, reducing data size per model via segmentation becomes beneficial; and (ii) distributing model-generated data across diverse models increases the model-dependent informativeness of the data each model receives. By contrast, increasing data per model (and reducing diversity) is only beneficial early on, when models are first exposed to new data and EDQ is dominated by sample size. The advantage of monoculture observed short-term can even disappear entirely under three modeling variations: introducing diversity by combining model families rather than segmenting data, instantiating diversity via dataset expansion rather than dataset segmentation, and evaluating community models on community-specific test sets rather than a universal benchmark. These findings, reported in the Appendix G, suggest that multiple forms of model diversity potentially can potentially be combined to mitigate collapse.
How should these results inform AI goals and governance? Most directly, the long-term epistemic risks of AI monoculture challenge the premise of a single universally optimal model, aligning with prior calls for diversity motivated by justice, fairness, and safety (Jain et al., 2024; Blili-Hamelin et al., 2025; Gebru and Torres, 2024; Lu and Kleek, 2024; Fisher et al., 2025; Lazar and Nelson, 2023). Importantly, promoting diversity in AI does not imply endorsing inferior models; rather, it acknowledges that there are multiple legitimate conceptions of what constitutes beneficial AI.
Concerns about AI-driven knowledge collapse (Peterson, 2025; Wright et al., 2025) reflect the need of evaluation frameworks that capture long-term ecosystem effects—effects that short-term benchmark performance may miss. Our results suggest that model similarity (i.e., ecosystem homogeneity) can serve as an indicator of collapse risk. We welcome recent efforts to measure epistemic diversity both within (Wright et al., 2025) and among (Zhang et al., 2025a) models, and encourage AI operators to assess their models’ disagreement relative to other models deployed in the same ecosystem.
Policy approaches from other domains may offer useful analogies, including biodiversity protection, DEI initiatives in science, and regulations that limit excessive genetic homogeneity (e.g., cousin-marriage laws). In AI, governance could reward model outputs increasing diversity in shared online data pools or regulate the distribution of critical resources such as training data. Diversity can also be incentivized bottom-up through model specificity: requiring systems to be explicit about the tasks, communities, languages, values, etc. they serve (Blili-Hamelin et al., 2025) would naturally push ecosystems toward greater differentiation. As long as models remain beneficial within their specified contexts, too much diversity should not be a concern.
Given the current trajectory toward AI monoculture (Lazar and Nelson, 2023; Sourati et al., 2025; Zhang et al., 2025a), a shift toward greater model diversity appears necessary to sustain long-term knowledge production.
Appendix A Model progress including performance on base models
@@
Figure 8 shows the perplexity progress of the two model families across ten iteration steps, including the models’ perplexity for the base models.
@@
Appendix B OPT-125m M=1 runs with different random seeds
Figure 9 shows the performance trajectories for $M=1$ across ten iterations for four independent runs. The similarity of these trajectories indicates that the results are largely independent of the random seed. Ecosystems with higher diversity already average performance across two or more models.
Figure 9. The performance trajectory of $M=1$ is largely run-independent. Mean perplexity and standard deviation for ecosystem $M=1$ for each of four independent runs with different random seeds.
Appendix C OPT-125m performance across 20 iterations
Figure 10 shows the optimal diversity level for OPT-125m across up to 20 self-training iterations ($T$), showing that the optimal diversity level continues to increase beyond 10 iterations.
(a) Optimal level of diversity.
(b) Performance progress.
Figure 10. The optimal level of ecosystem diversity continues to increase beyond 10 self-training iterations ($T$). (a) Optimal diversity level and the corresponding performance trajectories for OPT-125M. (a) and (b) extend the iteration range of Figures 3 and 4. Colored symbols correspond to the diversity levels shown in Figure 2 with an additional level of $D=8$ in light green.
Appendix D Illustrations of probability distributions and EDQ
Figure 11 illustrates the probability distributions in self-training loops, and Figure 12 illustrates the model-dependent EDQ.
Figure 11. Illustration of probability distributions in self-training loops. Because the training data is finite, the empirical distribution, $\hat{p}_{data(t=0)}$, diverges from the true distribution, ${p}_{data}$. The model’s inherent biases (functional expressivity and functional approximation errors) further distort the distribution it represents, which is $p_{\theta(t=0)}$, and the distribution of the data it generates, which is $\hat{p}_{data(t=1)}$.
Figure 12. Illustration of model-dependent EDQ. For the illustrated distribution discrepancies, the training data (“DATA”) is beneficial for “m2” (high EDQ) but harmful for “m1” (low EDQ). EDQ is defined as $\delta_{EDQ}(\hat{p}{data},p{\theta}):=\mathbb{W}2(\hat{p}{data(t)},p_{data})-\mathbb{W}2(p{\theta(t)},p_{data}),$ where lower values of $\delta_{EDQ}$ indicate higher EDQ. It is important to note that the introduced scalar metric serves as a heuristic to explain model-dependent data quality; therefore, we ignore additional factors shaping model weights.
Appendix E Perplexity distributions for both model families
Figure 13 shows the mixed perplexity distributions of the OPT-125m and GPT2 ecosystems on the original training dataset, indicating that model collections with greater diversity exhibit both greater variety in the information represented and lower alignment with the training data distribution.
(a) OPT-125m
(b) GPT2
Figure 13. Perplexity distributions widen with increasing diversity. Mixed perplexity distribution on the original Wikitext2 train data points, evaluated by all models in a given ecosystem with diversity ($D$ =$M$ = 1, 2, 4, 16) at iteration steps $t$ = 0 (left) and $t$ = 9 (right). As number of models increases, the distributions become wider, indicating greater diversity. Simultaneously, their means shift toward higher perplexity, suggesting greater divergence from the distribution in the original training set. Across all ecosystems, the distributions narrow from the beginning to the end of the model iterations.
Appendix F GPT2 V1: Model size and data size
Figure 14 shows GPT2 performance results for increased model size (GPT2-Medium) and larger data size (2.12.1M tokens).
(a) Optimal diversity levels.
(b) Aggregated perplexities after 10 iterations.
(c) Mean perplexities for M;D=1
(d) Mean perplexities for M;D=16
[…]
Appendix G Additional experimental variations of the primary setting
The observed short-term benefit of homogeneous ecosystems is due to the segmentation approach, which reduces the data size per model as diversity increases, and to the use of a universal ground-truth dataset (the Wikitext2 test set in our case). This initial advantage disappears when diversity is instantiated through combining two model families (Figure 16), through data expansion (Figure 17(b)), or when community-specific test sets are employed. The aggregated results are shown in Figure 15.
Diversity through combining 2 model families
Real-world diversity may arise from multiple system-design dimensions, including pre-training data, architecture, alignment, interface design, and system messaging. Figure 16 shows that combining two model families (OPT-125m and GPT2) consistently outperforms the combination of two models of same family, suggesting that diversity across design dimensions may further amplify benefits.
Diversity via data expansion
The expansion approach reflects a more realistic landscape of multiple similar models (Lazar and Nelson, 2023; Zhang et al., 2025a). If dominant models are trained on largely overlapping internet-scale data (Villalobos et al., 2024; Mozilla, 2024), segmentation through reduction may be the only feasible diversification strategy; otherwise, diversity could be achieved via expansion by incorporating models trained on previously unused data. In such comparisons, diverse ecosystems consistently achieve lower perplexity across iterations (Figure 17(b)), whereas simply adding more similar models does not prevent collapse (Figure 20(a)).
Community-specific test sets
Evaluating community models using community-specific datasets rather than universal benchmarks yields a more realistic assessment; under this setting, diverse ecosystems outperform homogeneous ones already at $t=0$ (Figure 17(a)). Community- or domain-specific models challenge the prevailing assumption that AI systems should serve everyone equally across all tasks (Blili-Hamelin et al., 2025; Bender et al., 2021). Granting affected communities greater agency in AI design (Costanza-Chock, 2020) may naturally guide ecosystems toward healthier diversity.
Human language and knowledge evolve alongside AI systems. Following prior work (e.g., Herel and Mikolov, 2024; Shumailov et al., 2024; Alemohammad et al., 2023), we use a static dataset for generation and evaluation; in practice, this would resemble evaluating future models on today’s data. A more realistic approach would involve evolving generation and evaluation datasets over time.
While further research is needed to identify optimal ecosystem parameters under realistic conditions, our results provide strong evidence that diversity is a key factor in sustaining long-term AI ecosystem performance.
Figure 15. The advantage of homogeneous ecosystems observed short-term can even disappear entirely under three modeling variations: introducing diversity by combining model families rather than segmenting data (“OPT-125m(+GPT2)” and “GPT2(+OPT-125m)”) instantiating diversity via dataset expansion rather than dataset segmentation (“OPT-125m, Expansion”), and evaluating community models on community-specific test sets rather than a universal benchmark (“OPT-125m, Community”). These variations involved only two diversity levels ($D=1,2$).
(a) Combination of two OPT-125m models compared with an OPT-125m paired with GPT-2
(b) Combination of two GPT-2 models compared with an GPT-2 paired with OPT-125m
Figure 16. Two diverse pre-trained models yield lower perplexity than two identical pre-trained models. OPT-125m+OPT-125m and GPT2+GPT2 represent ecosystems composed of two identical models (green). OPT-125m(+GPT2) and GPT2(+OPT-125m) represent ecosystems where each model is paired with the other type (orange). Unlike the primary segmentation approach, diversity here is instantiated by using two different pre-trained models. All models are trained on the same dataset at iteration step $t=0$ and use the same generation dataset throughout; thus, aside from their pre-training differences, the training procedure is identical across models. The shaded region represents the combined standard deviation across models and data points.
(a) From the first iteration step ($t=0$) onward, two community models consistently achieve lower perplexity than a single model trained and evaluated on the combined datasets. $M=1$: a single model trained and evaluated on the combination of Wikitext2 and Penn Treebank (green). $M=2$: two community models, each trained and evaluated on its corresponding dataset (orange). To ensure a controlled comparison, the length of the Wikitext2 dataset is reduced to match the length of Penn Treebank.
(b) Two models trained on two different segments of Wikitext2 consistently achieve lower perplexity than two models both trained on one of the two segments. $D=1$: Two models trained on identical segment 0 (green). $D=2$: Two models trained on two different segments 0 and 1 (orange), corresponding to the same ecosystem as M=2 in the primary setting. Unlike in the primary segmentation approach, diversity is achieved through expansion of the distribution.
Figure 17. The shaded region is a combined standard deviation across the models and data points.
Appendix H Sensitivity analysis
We conducted additional experiments to isolate factors that might affect the performance changes observed in the primary setting. The results indicate that neither (model-independent) data quality (Figure 18), training data size (Figure 19), the number of models per ecosystem (Figure 20), nor overfitting (Figure 21) can fully account for the observed performance decay or the differences across ecosystems. However, overfitting, which is captured by the EDQ, appears to be an important contributor to performance decay.
Because these individual factors fail to explain the performance trajectories in isolation, the findings provide support for the EDQ hypothesis.
(a) OPT-125m separately trained on generated output
(b) OPT-125m tested on generated output
Figure 18. Data quality alone cannot explain performance differences across ecosystems. The OPT-125m base model was separately trained (a) and tested (b) on generated outputs from the least and most diverse ecosystems ($D=1$ and $D=16$, respectively). For $D=16$, outputs from all 16 models were aggregated. In (a), the off-the-shelf OPT-125m base model was fine-tuned separately on each generated dataset and subsequently evaluated on the fixed Wikitext2 test set. The similar perplexity values obtained for $D=1$ and $D=16$ indicate that model-independent data quality cannot fully account for the performance differences observed in our main experiments. Lower perplexity values for outputs from iterations $t=0$ and $t=1$ in (b) suggest that artificial data generated at early stages of the self-training loop more closely resembles the pretraining distribution of the base OPT-125m model. Fine-tuning on such artificial data shifts the distribution of the fine-tuned model back toward the pretraining distribution and away from the true distribution, which helps explain the performance drop observed with the output from iteration $t=0$ in (a) and consistently across our experiments.
(a) Single OPT-125m re-trained on the real Wikitext2 train set
(b) Single OPT-125m self-trained on own output
Figure 19. Training data size alone cannot explain performance differences across ecosystems. A single OPT-125m model was analyzed in two conditions: (a) re-trained on the real Wikitext2 training set and (b) self-trained on its own output, respectively. The initial training data content and sizes correspond to the primary setting (N=0.8N=0.8M tokens), but only one model was used in each run. The condition “$D=1$, OPT-125m, NN” matches $D=1$ in the primary setting. The similar perplexity values observed across different data sizes indicate that training data size alone cannot account for the performance differences across ecosystems.
(a) OPT-125m
(b) GPT-2
Figure 20. Number of models cannot fully explain the performance differences observed across ecosystems. 1×OPT-125m and 1×GPT-2: a single model trained on the entire dataset of size $N$, corresponding to the $D=1$ condition in the primary setting (green). 2×OPT-125m and 2×GPT-2: two identical models, each trained on the entire dataset of size NN (orange). The generation datasets are identical for all models. The training datasets are identical at the first iteration step but may differ slightly in later iterations for the two-model runs due to stochasticity in setup. The shaded region represents the combined standard deviation across models and data points.
(a) OPT-125m
(b) GPT-2
(c) OPT-125m evaluated on train set
(d) GPT-2 evaluated on train set
Figure 21. Overfitting alone cannot explain the observed performance decay. Comparison of the self-training loop in the primary setting (green) with a re-training loop on the fixed real training set (orange), evaluated on the test set (top) and training set (bottom). Because the training data in the re-training condition remains fixed, any performance decay can be fully attributed to overfitting. The lower perplexity on the training set in the re-training condition indicates stronger (over)fitting to the fixed training data. This suggests that overfitting alone cannot account for the performance decline observed in the self-training loop, which appears comparable for OPT-125m and even stronger for GPT-2.
Appendix I Hyperparameters
Table 1 provides the hyperparameters used for text generation.
Hyperparameter Value
temperature 0.0
do_sample False
max_new_tokens 128
min_new_tokens 128
repetition_penalty 0.0
batch size 32
block size 128
num_beams 5
Table 1. Generation hyperparameters.